OpenStudy (anonymous):

4) Find all values where the function is discontinuous. f(x) = 4x/((3x-1)(2x+5))

OpenStudy (anonymous):

Btw this is just review for my midterm on my week points for calculus.

OpenStudy (amistre64):

there are relatively few cardinal sins in math; the biggest is dividing by a zero

OpenStudy (anonymous):

So i really appreciate the help, I have been so busy with my senior project and my other class that I have had very little time to study.

OpenStudy (amistre64):

what values of x make the bottom go zero?

OpenStudy (anonymous):

Finding the zero's?

OpenStudy (amistre64):

the zeros of the bottom, not the zeros of the function perse

OpenStudy (anonymous):

I remember trying to solve for zero

OpenStudy (amistre64):

then solve for zeros of the bottom: (3x-1)(2x+5) = 0, for what values of x?

OpenStudy (anonymous):

making the equation equal to zero?

OpenStudy (anonymous):

yes exactly

OpenStudy (amistre64):

recall your multiplication tables .... 0 was one of the simplest ones to remember

OpenStudy (amistre64):

given a product of a and b ab = 0 when a=0, or b=0, or both equal 0

OpenStudy (anonymous):

Anything times zero is zero lol

OpenStudy (amistre64):

yes :) so (3x-1)(2x+5) = 0, when 3x-1 = 0 , or when 2x+5 = 0

OpenStudy (anonymous):

4x/((3x-1)(2x+5)) 4x = 0 3x -1 = 0 2x + 5 = 0

OpenStudy (amistre64):

4x=0 is not a cardinal sin in this case. we are totally allowed to have a zero on top of a fraction

OpenStudy (amistre64):

a function is discontinuous where it is undefined, or where it breaks apart

OpenStudy (amistre64):

a fraction is undefined when it has a zero underneath: n/0 is bad

OpenStudy (anonymous):

I get that

OpenStudy (amistre64):

so we see that this function is undefined when: 3x -1 = 0 or 2x + 5 = 0 so all we have to do is determine for what values of x that happens

OpenStudy (anonymous):

So the numerator is allow to have a zero but the denominator is not allowed

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

So fill in values for x?

OpenStudy (amistre64):

of course, or algebra out some values

OpenStudy (amistre64):

im not real sure what process youre thinking of there

OpenStudy (amistre64):

3x - 1 = 0 , to solve for x, lets start by adding 1 to each side + 1 +1 ---------- 3x + 0 = 1 , since anything +0 is itself (indentity), this simplifies to 3x = 1 , now to get rid of the 3 stuck there, we divide it off, 3/3 = 1 /3 /3 ------- 1x = 1/3 , since anything times 1 is itself (identity), this simplifies to x = 1/3 the same concepts can be applied to the other factor

OpenStudy (anonymous):

3x -1 = 0 3x = 0 + 1 or 2x + 5 = 0 2x = 0 + 5

OpenStudy (amistre64):

2x = -5 , but yeah

OpenStudy (anonymous):

yeah i forgot to flip the sign

OpenStudy (amistre64):

if your intent is to "flip the sign" and "move things to the other side"; then you are not applying any sound mathematical principals to this

OpenStudy (amistre64):

there are 5 simple properties to algebra, and they can prevent a world of mistakes

OpenStudy (amistre64):

1) ab = ba , commutative property 2) (ab)c = a(bc) , associative property 3) \(aa^{-1}=e\), inverse property; the inverse in addition is the negative, the inverse in multiplication is dividing by the reciprocal 4) \(ae=a\), identity property; the identity in addition is +0, the identity in multiplication is *1 5) a(b+c) =ab + ac , distributive property ... multiplication distributes over addition

OpenStudy (amistre64):

2x + 5 = 0 , addition inverse, -5 - 5 -5 ----------- 2x + 0 = -5 , addition identity 2x = -5 , multiplicative inverse, /2 /2 /2 -------- 1x = -5/2 , multiplicative identity x = -5/2 , multiplicative identity

OpenStudy (anonymous):

okay so then we try to get x alone

OpenStudy (anonymous):

I understand that

OpenStudy (amistre64):

yes, by using the properties of algebra, we can isolate the variable

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

signs reverse on the other side of the equation

OpenStudy (amistre64):

:) or make up some properties as we go .... the sign reverse on the other side property for example. head math is great, but it has its drawbacks is all

OpenStudy (anonymous):

okay so the inverse becomes the identity

OpenStudy (amistre64):

correct, applying the inverse of a given element, produces the identity element

OpenStudy (anonymous):

inverse meaning the opposite and identity the after math

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

so this function is discontinuous at x = -5/2

OpenStudy (amistre64):

that is one of the places yes

OpenStudy (amistre64):

the other one is at x=1/3

OpenStudy (amistre64):

the strategy in determining discontinuity is in finding the parts of the function that can "go bad" dividing by 0 is bad taking the even-root of a negative number is bad among the Reals taking the logarothm of 0 or less is bad those seem to be the big ones to watch for

OpenStudy (anonymous):

Okay thanks you up for another one with lim?

OpenStudy (amistre64):

i can take a stab at it .. i only have so much wits about me in a day before I start to go stupid lol

OpenStudy (anonymous):

lmao yeah thats how I feel right about now